Station Messages
21 Jan 2020 20:08 UTC 2020021+20:08 UTC

Overview Datums Bench Marks Data Query Site Visit Daily QC Messages Attributes

134: Bahia Grande Water Quality Station #3

26° 1' 11" N   97° 16' 28" W
BGWQ3
University of Texas at Brownsville and Texas Southmost College University of Texas at Brownsville and Texas Southmost College

2010306+1509 (acknowledged, lloyd)

No data received since 2010306+0000.

2010216+1108 (retired, ussery)

No data since 2010118+2100.

2010095+1049 (retired, ussery)

Problems with pH and conductivity probes.

2009354+1159 (retired, jrizzo)

Battery voltage at 1 V. Battery up, data resumed.

2009310+1224 (retired, ussery)

No data received since 2009230+0130.

2008331+1557 (retired, ussery)

No data since 2008201.

2008233+1642 (retired, ussery)

No data since 2008230. No response from tower.

2008210+1309 (retired, ussery)

No data since 2008206+1700. Power outage at Los Fresnos.

2008199+1135 (retired, ussery)

No data since 2008188+1400.

2008185+1329 (retired, ussery)

No data since 2008183+2200. Tower struck by lightning.

2008113+2011 (retired, ussery)

No data since 2008112+1830. Data has resumed.

2008053+1307 (retired, ussery)

No data since 2008038+1630. Data has resumed.

2008036+1205 (retired, jrizzo)

DO, PH suspect.

2007317+1359 (retired, ussery)

Sporadic data received 2007302+0700.

2007285+1349 (retired, rperez)

No data since 2007279. Data resumed. JD fix problem in poling program.

2007249+1217 (retired, jrizzo)

Specific conductance at zero. SPCON resumed.

2007219+1313 (retired, ussery)

No data since 2007200+0600. Service trip planned week of Aug. 6th.

2007194+2019 (retired, ussery)

No data since 2007186+0600. Data resumed after field crew visit.

2007093+1133 (retired, jrizzo)

No data since 2007091. Data resumed.

2007079+2019 (retired, jdavis)

No data since 071 -- tower pc replaced, data resumed

2006349+1213 (retired, jrizzo)

Hydrolab removed from poll because it was sending repeated text on connect:
Hydrolab H20 V1.02
©opyright 1990 Hydrolab Corporation. Data resumed

2006248+1316 (retired, jadams)
bat : today (01/21/2020 00:00 - 01/22/2020 00:00) :

Solar panels cleaned by UTB.

2006227+2216 (retired, jdavis)

Sporadic data since 2006226+1700.

2006221+1313 (retired, ussery)

No data imported since 2006219+1900. Data is imported.

2006209+1149 (retired, jrizzo)

Battery not charging. Solar panel covered in bird droppings - needs cleaning. Charging resumed

2006034+1800 (retired, jdavis)

No data since 023Los Fresnos tower down.
PC back up is collecting data. Satellite modem is out. Data is being downloaded automatically over telephone modem every 8 hours until satellite service is restored.
----
Satellite modem is back up and data is being transferred.

2005218+1408 (retired, kjenkin)

No data since 2005215.